# https://sinhasaptarshi.github.io/escounts

### v1

@@ 5 Conclusion
We have proposed to utilise exemplars for video repetition counting. We introduce Every Shot Counts (ESCounts), an attention-based encoder-decoder that learns to correspond exemplar repetitions across a full video. We define a learnable zero-shot latent that learns representations of generic repetitions, to use during inference. Extensive evaluation on RepCount, Countix, and UCFRep demonstrates the merits of ESCounts achieving state-of-the-art results on the traditional MAE and OBO metrics and the newly introduced RMSE and OBZ. Further analysis showed that the exemplars do not need to be from the same video as long as they belong to the same action class. The diversity of these exemplars is an aspect for future exploration.
Acknowledgements
This work uses publicly available datasets and annotations to produce all results and ablations. Research is supported by EPSRC UMPIRE (EP/T004991/1). S. Sinha is supported by EPSRC DTP (Doctoral Training Program) studentship.
@@ Acknowledgements
Code is made available anonymously from: https://github.com/sinhasaptarshi/EScounts. We provide the full code, checkpoint as well as a demo with a couple of videos to try our inference.
In this supplementary, we explain the attached qualitative video in Sec. 6. We then provide additional ablations on the architecture’s choices (e.g. depth of transformer and window size) in Sec. 7. Additionally, we evaluate the ability of ESCounts to locate each repetition within the video in Sec. 8. Finally, we analyse our improvement over baseline by studying groups of videos based on their length, number of repetitions, and counts in Sec. 9.
6 Qualitative Video and Extended Figure
We attach a compilation video (1380.mp4) showcasing our method’s abilities in Video Repetition Counting (VRC). For a diverse set of 11 videos from all 3 datasets, we show the video alongside synchronised ground truth and ESCounts predicted density maps. For each example, we note which test set it is from.
Additionally, we extend Fig. 5 in the main paper with additional examples in Fig. 8.
7 Further Ablations
We extend the ablation studies in  Sec. 4.3 by reporting results over different $L$ and $L^{\prime}$. Additionally, we analyse the impact of windowed-self attention on the performance of ESCounts.
(o) RepCount
(p) Countix
Figure 7: Additional qualitative results.
(a) UCFRep
Figure 8: Additional qualitative results (continued).
@@ 7 Further Ablations
Impact of $L$. We ablate the value of $L$ i.e. the number of layers in the cross-attention block. Increasing $L$ increases the number of operations that discover correspondences between the video and the selected exemplars. As seen in Tab. 7, while low $L$ causes a drop in performance, high $L$ can also be detrimental probably due to overfitting. $L=2$ gives the best results for the majority of the metrics.
Next keeping $L=2$ fixed, we vary $L^{\prime}$ in Tab. 7. $L^{\prime}$ is the number of windowed self-attention layers in the self-attention block. $L^{\prime}=3$ gives the best results on all the metrics. Similarly, increasing or decreasing $L^{\prime}$ drops performance gradually.
Self-attention vs Windowed Self-attention. Motivated by [35], we use windowed self-attention for the self-attention block in ESCounts. Given spatio-temporal tokens $\mathcal{T}^{\prime}\times H^{\prime}\times W^{\prime}\times C$, windowed self-attention computes multi-headed attention for each token within the immediate neighbourhood using 3D shifted windows of size $t^{\prime}\times h^{\prime}\times w^{\prime}$, where $t^{\prime}<\mathcal{T}^{\prime}$, $h^{\prime}<H^{\prime}$ and $w^{\prime}<W^{\prime}$. Here we ablate for various values $(t^{\prime},h^{\prime},w^{\prime})$ in Tab. 7. Note that $t^{\prime}=\mathcal{T}^{\prime}$, $h^{\prime}=H^{\prime}$ and $w^{\prime}=W^{\prime}$ is using standard self-attention over all the embeddings, where each token is attended with every token and is denoted by full. As shown, the best performance is obtained with window size $(4,7,7)$, demonstrating the importance of attending to tokens in immediate spatio-temporal neighbourhoods only. We found that the results are most sensitive to the value of $t^{\prime}$. Performance steadily drops as the value of $t^{\prime}$ changes.
Sampling Rate for Encoding. As stated in the implementation details, we construct our input by sampling every four frames from the video. We ablate the impact of the sampling rate in Tab. 8. From the results, we observe that denser sampling is key for robust video repetition counting. Reducing the sampling rate steadily decreases performance.
@@
Table 9: OBO, parameters, and training and inference speeds on UCFRep. Mtreics obtained by the public available codebase of [65] are denoted with .
@@
Model Size and Speed For UCFRep [65], [65, 67] achieve better performance than ESCounts. However, this performance is achieved by having more trainable parameters, as [65, 67] finetune the encoders on the target dataset. We use the provided codebase from [65] and benchmark the average number of iterations per second for a full forward and backward pass over the entire training set alongside inference-only average times on the test set. We use the same experiment set-up described in Sec. 4.1 and report speeds in Table 9. Training ESCounts is  8 times faster. Interestingly, even during inference ESCounts is  12 times faster than Context [65] which uses iterative processing. Note that [67] could not be used for this analysis as their code for training with UCFRep is not publicly available.
@@ 8 Repetition Localisation
Figure 9: Explaining localisation metric $\mathcal{J}$. We identify local maxima in $\tilde{\mathbf{d}}$ and threshold peaks higher than $r$ to remove noise. $\mathcal{J}$ is then computed between the annotated start-end times and the thresholded peaks.
@@
VRC metrics only relate predicted to correct counts, regardless of whether the repetitions have been correctly identified. We thus investigate whether the peaks of the predicted density map $\tilde{\mathbf{d}}$ align with the annotated start-end times of repetitions in the ground truth. Following action localisation methods [8, 18, 22], we adopt and adjust the Jaccard index $\mathcal{J}$ for repetition localisation. Using the predicted density map we define different thresholds for salient regions of repetitions. As the values of $\tilde{\mathbf{d}}$ peaks vary across videos, we apply thresholds $\theta$ relative to the maximum and minimum values, $r=\theta(\max(\tilde{\mathbf{d}})-\min(\tilde{\mathbf{d}}))$. We find all local maxima in $\tilde{\mathbf{d}}$ and only keep those above threshold $r$. We consider a repetition to be correctly located (TP) if at least one peak occurs within the start-end time of that repetition. Peaks that occur within the same repetition are counted as one. In contrast, peaks that do not overlap with repetitions are false positives (FP) and repetitions that do not overlap with any peak are false negatives (FN). We then calculate the Jaccard index as TP divided by all the correspondences (TP + FP + FN) as customary.
(a) counts
(b) repetition length
(c) video duration
Figure 10: Comparison of ESCounts and TransRAC [20] for groups of counts/repetition duration/video-length
@@
In Table 10 we report the Jaccard index over different thresholds alongside the Mean Average Precision (mAP) on RepCount. We select TransRAC [20] as a baseline due to their publicly available checkpoint. Across thresholds, ESCounts outperforms [20] with the most notable improvements observed over higher threshold values demonstrating that ESCounts predict density maps with higher contrast between higher and lower salient regions. For 0.9, 0.8, and 0.7 thresholds ESCounts demonstrates an +8.68%, +10.30%, and +8.10% improvement over [20].
9 Comparison with TransRAC over Groups
As in Sec. 4.3, we divide the videos into groups based on their counts, average repetition duration, and lengths. We compare ESCounts’ performance against TransRAC in each of these groups. As evident from Fig. 10, ESCounts performs significantly better than TransRAC on high counts. From Fig. 10, ESCounts performs strongly over varying repetition duration, whilst TransRAC struggles on videos with XS and XL repetitions. Fig. 10 shows that even though both TransRAC and ESCounts performance drops on longer videos, ESCounts perform significantly better on them.

### v2

@@ 5 Conclusion
We have proposed to utilise exemplars for video repetition counting. We introduce Every Shot Counts (ESCounts), an attention-based encoder-decoder that learns to correspond exemplar repetitions across a full video. We define a learnable zero-shot latent that learns representations of generic repetitions, to use during inference. Extensive evaluation on RepCount, Countix, and UCFRep demonstrates the merits of ESCounts achieving state-of-the-art results on the traditional MAE and OBO metrics and the newly introduced RMSE and OBZ. We provide detailed analysis and ablations of our method, highlighting the importance of training with exemplars and time-shift augmentations. The diversity of these exemplars is an aspect for future exploration.
Acknowledgements. This work uses publicly available datasets and annotations for results and ablations. Research is supported by EPSRC UMPIRE (EP/T004991/1). S. Sinha is supported by EPSRC DTP studentship.
@@
Code is made publicly available at: https://github.com/sinhasaptarshi/EveryShotCounts. The repository contains the full train and evaluation code and a demo for inference with a few videos.
In the following sections, we provide more qualitative results in Sec. 6. We then provide additional ablations on the architecture’s choices (e.g. depth of transformer and window size) in Sec. 7. Additionally, we evaluate the ability of ESCounts to locate each repetition within the video in Sec. 8. We then compare VRC to Temporal Action Segmentation (TAS) in Sec. 9 demonstrating distinctions between the two tasks.
Additionally, following the release of the recent egocentric video counting dataset OVR-Ego4D [16], we train and evaluate ESCounts on this newly introduced dataset demonstrating the effectiveness of our method for egocentric counting in Sec. 10.
@@
6 Qualitative Video and Extended Figure
We provide a compilation of videos on our website https://sinhasaptarshi.github.io/escounts/ showcasing our method’s Video Repetition Counting (VRC) abilities over a diverse set of 20 videos from all 3 datasets. Videos are shown alongside synchronised ground truth and predicted density maps. The test set from which each video is sampled is also shown.
We additionally extend Fig. 5 in the main paper with more examples from all datasets in Fig. 8.
7 Further Ablations
We extend the ablations in Sec. 4.3, report results over different $L$ and $L^{\prime}$, and analyse the impact of windowed-self attention on the performance of ESCounts.
(o) RepCount
(p) Countix
Figure 7: Additional qualitative results.
(a) UCFRep
Figure 8: Additional qualitative results (continued).
Impact of $L$. We ablate $L$ i.e. the number of layers in the cross-attention block. Increasing $L$ increases the number of operations that discover correspondences between the video and the selected exemplars. As seen in Tab. 7, while low $L$ causes a drop in performance, high $L$ can also be detrimental probably due to overfitting. $L=2$ gives the best results for the majority of the metrics.
Next keeping $L=2$ fixed, we vary $L^{\prime}$ in Tab. 7. $L^{\prime}$ is the number of windowed self-attention layers in the self-attention block. $L^{\prime}=3$ gives the best results across all the metrics. Similarly, increasing or decreasing $L^{\prime}$ drops performance gradually.
Self-attention vs Windowed Self-attention. Motivated by [40], we use windowed self-attention for the decoder self-attention blocks. Given spatio-temporal tokens $\mathcal{T}^{\prime}\times H^{\prime}\times W^{\prime}\times C$, windowed self-attention computes multi-headed attention for each token within the immediate neighbourhood using 3D shifted windows of size $t^{\prime}\times h^{\prime}\times w^{\prime}$, where $t^{\prime}\leq\mathcal{T}^{\prime}$, $h^{\prime}\leq H^{\prime}$ and $w^{\prime}\leq W^{\prime}$. We ablate on various $(t^{\prime},h^{\prime},w^{\prime})$ values in Tab. 7. Note that for $t^{\prime}=\mathcal{T}^{\prime}$, $h^{\prime}=H^{\prime}$, and $w^{\prime}=W^{\prime}$ denoted as full, standard self-attention is used where each token attends to every token. As shown, the best performance is obtained with window size $(4,7,7)$, demonstrating the importance of attending to tokens in immediate spatio-temporal neighbourhoods only. We found variations in the value of $t^{\prime}$ to have the largest performance impact with decreases as the value of $t^{\prime}$ changes.
Sampling Rate for Encoding. As stated in the implementation details, we sample every four frames from the video to form the encoder inputs. We ablate the impact of the sampling rate in Tab. 8. As shown, denser sampling is key for robust video repetition counting. Reducing the sampling rate steadily decreases performance as relevant parts of repetitions may be missed.
@@ 7 Further Ablations
Table 9: OBO, parameters, and training and inference speeds on UCFRep. Metrics obtained by the public available codebase of [70] are denoted with .
@@
Model Size and Speed For UCFRep [70], [70, 72] achieve better performance than ESCounts. However, this performance is achieved by having more trainable parameters, as [70, 72] finetune the encoders on the target dataset. We use the provided codebase from [70] and benchmark the average number of iterations per second for a full forward and backward pass over the entire training set. Additionally, we report inference-only average times on the test set. We use the same experiment set-up described in Sec. 4.1 and report speeds in Tab. 9. Training ESCounts is $\sim$8$\times$ faster. Interestingly, ESCounts maintains its efficiency even during inference with $\sim$12$\times$ faster times than Context [70] which uses iterative processing. Note that [72] could not be used for this analysis as their code for training with UCFRep is not publicly available.
@@
Figure 9: Localisation metric $\mathcal{J}$. We identify local maxima in $\tilde{\mathbf{d}}$ and threshold peaks higher than $r$ to remove noise. $\mathcal{J}$ is then computed between the annotated start-end times and the thresholded peaks.
@@ 8 Repetition Localisation
VRC metrics only relate predicted to correct counts, regardless of whether the repetitions have been correctly identified. We thus investigate whether the peaks of the predicted density map $\tilde{\mathbf{d}}$ align with the annotated start-end times of repetitions in the ground truth. Following action localisation methods [8, 20, 25], we adopt the Jaccard index $\mathcal{J}$ for repetition localisation. As the values of $\tilde{\mathbf{d}}$ peaks vary across videos, we apply thresholds $\theta$ relative to the maximum and minimum values, $r=\theta(\max(\tilde{\mathbf{d}})-\min(\tilde{\mathbf{d}}))$. We find all local maxima in $\tilde{\mathbf{d}}$ and only keep those above threshold $r$. We consider a repetition to be correctly located (TP) if at least one peak occurs within the start-end time of that repetition. Peaks that occur within the same repetition are counted as one. In contrast, peaks that do not overlap with repetitions are false positives (FP) and repetitions that do not overlap with any peak are false negatives (FN). We then calculate $\mathcal{J}$ as TP divided by all the correspondences (TP + FP + FN) as customary.
@@
In Tab. 10 we report the Jaccard index over different thresholds alongside the Mean Average Precision (mAP) on RepCount. We select TransRAC [22] as a baseline due to their publicly available checkpoint. Across thresholds, ESCounts outperforms [22] with the most notable improvements observed over higher threshold values. This demonstrates ESCounts’ ability to predict density maps with higher contrast between higher and lower salient regions. For 0.9, 0.8, and 0.7 thresholds ESCounts demonstrates a +8.68%, +10.30%, and +8.10% improvement over [22].
Table 11: Comparison between ESCounts and TAS baseline on close and open-set RepCount setting.
Task Method benchmark open-set
[…]
TAS GTRM [24] 0.527 0.159 1.000 0.000
TriDet [53] 0.603 0.232 1.000 0.000
VRC ESCounts 0.213 0.563 0.436 0.519
9 Distinction between VRC and TAS
Unlike Temporal Action Segmentation (TAS) methods, VRC methods can generalise to unseen action classes. In Tab. 11 we compare ESCounts to a TAS method [24] on the RepCount benchmark (close-set) and open-set setting. As shown, [24] can only localise the actions of a pre-defined set of categories with which the model was trained. In contrast, VRC is learned as an open-set task. As ESCounts uses a learnt latent to encode class-independent repetition embeddings, it effectively generalises to unseen categories. In addition, ESCounts can better handle large variations in repetition durations that are present in VRC videos compared to [24], which as noted by [22] is a weakness of TAS methods.
10 Results on egocentric VRC.
Table 12: Results on OVR-Ego4D.$\dagger$ indicates results have been copied from [16]. (V) corresponds to vision-only models and (V+L) to vision and language models.
[…]
V RepNet [15] $\dagger$ 3.20 0.74 0.19 0.43
ESCounts 2.41 0.32 0.30 0.68
V+L OVRCounter [16] $\dagger$ 1.60 0.35 0.29 0.66
The recently-introduced OVR-Ego4D [16] is an Ego4D [19] subset containing clips of repetitive egocentric actions, e.g. cutting onions, rolling dough. It comprises 50.6K 10-second clips with 41.9K train and 8.7K test clips. Annotations are only provided for the number of repetitions and not the individual start and end times per repetition. Thus, similar to Countix, we define pseudo-labels to estimate the density maps.
We evaluate ESCounts on OVR-Ego4D in Tab. 12. Compared to the vision-language-based OVRCounter, [16] ESCounts improves OBZ, OBO, and MAE, with only visual inputs, without any language input in training or inference, showing ESCounts’ effectiveness for the domain of egocentric counting. We also add some qualitative results in Fig. 10. Similar to results on other datasets, ESCounts predicts accurate counts a over diverse range of counts. The peaks of individual repetitions are not as clear, due to the pseudo-labels, but ESCounts correctly finds the OBO counts in each case.
Figure 10: Qualitative results of ESCounts on OVR-Ego4D. For the selected videos, we show both ground truth (GT) and predicted (P) density maps along with the counts. Note that for OVR-Ego4D, we do not have temporal annotations for individual repetitions. Therefore similar to Countix, we show pseudo-labels as the GT density maps.